OpenStudy (anonymous):

find the slope of tangent line to the given curve at t=-2

OpenStudy (abb0t):

You're missing some iforamtion.

OpenStudy (anonymous):

Here you go

OpenStudy (abb0t):

first step is to find the derivative(s). Can you do that?

OpenStudy (abb0t):

That way, you can get the slope of the tangent line.

OpenStudy (abb0t):

Note that: \(\huge \frac{ dy }{ dx } = \frac{ \frac{ dy }{ dt } }{ \frac{ dx }{ dt } } \)

OpenStudy (anonymous):

;) sorry because I was in a hurry I didnt pay attention